The lightning rod was hooked into the grid by maintenance in the last week. We also ordered a new probe from Campbell and then went on a camera search for a camera that will actually work as opposed to our current option.
On the roof we discovered a small amount of water inside the East side box. We need to seal the boxes better with ‘monkey snot’. We had a compass to adjust the magnetic north for the weather station from the lightning rod installation that shifted it a bit.